# On the number of modes of Gaussian kernel density estimators

### v2

@@ 3.1 Bounding the third order error
$\eta_{3}\coloneqq\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$. Now, by a similar method as Lemma 2.3, we obtain the following bound. It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
@@
Recall $T$ from 1.2 and $A_{t}$ from Lemma 2.3. Then
$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
Proof of Lemma 3.2.
By the proof of Lemma 2.3, 3.2, and 3.1
$\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$=\frac{1}{6}\sum_{k=0}^{3}\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\kappa_{t}^{(k,3-k)}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left[\varphi H^{(k,3-k)}}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\asymp\int_{T}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}\eta_{3}e^{-A_{t}}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}y}\right)^{O(1)}e^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\,\mathrm{d}y\,\mathrm{d}t$
[…]
@@
where the last step follows 2.11 and the definition 1.2 of $T$. ∎
@@
One can see at this is actually an asymptotic equality by checking the Gaussian integrals in the proof above are of their typical order (i.e. no cancellation of leading terms). Hence, the decay is only a factor of $e^{-\omega(\beta)/4}$. For $t\not\in T$, even $t=\sqrt{2\log n-0.99\log\beta}$, the last inequality in Lemma 3.2 fails and we get a bound of polynomially larger that $\sqrt{\beta}$. Then, as the third order error is asymptotically larger than the contribution of the Gaussian approximation, so the normal approximation is no longer valid. This can be seen by comparing the plots in Figure 4.
@@ 3.2 Bounding higher order errors
We follow the classical proof about the validity of the Edgeworth expansion as an asymptotic series to show bound the higher order pointwise error of density function as follows.
@@
Assume that $p_{t}$ is bounded almost everywhere for any fixed $t\in T$. If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, then for any $t\in T$ and $\mathbf{x}\in\mathbb{R}^{2}$, we have that
$\mathopen{}\mathclose{{}\left(1+\|\mathbf{x}\|^{3}}\right)\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim e^{-\frac{\omega(\beta)}{4}}.$ (3.3)
We defer the proof to Section A.3. Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$. There, it is shown that the left hand side is $o(n^{-\frac{1}{2}})$ provided third moments $\eta_{3}=O(1)$. With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-\frac{1}{2}}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
@@
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for $c>0$, then asymptotically in $n,\beta\to\infty$
$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}.$
Proof of Corollary 3.5.
By Lemma 3.4 and computations in the proof of Lemma 2.3
@@
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\mathopen{}\mathclose{{}\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right\|^{3}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left(1+\Theta\mathopen{}\mathclose{{}\left(\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}}\right)^{\frac{3}{2}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}}\right)^{-2}\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}$
$\asymp e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
for $\tilde{y}\asymp\beta^{1/4}e^{{t^{2}}/{4}}y$, where we note that $T$ has length $\Theta(\sqrt{\log\beta})$ by 1.2 and
$\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{3}}\,\mathrm{d}\tilde{y}=\frac{2\pi}{3\sqrt{3}}=O(1).\qed$
4 Proof of Theorem 1.2
We prove Propositions 1.6 and 1.7 by checking 2.3, thereby proving Theorem 1.2.
4.1 Proof of Proposition 1.6
To prove Proposition 1.6 we seek to apply Theorem 2.1 to $\mathbb{E}U_{0}(F_{n},T)$. This in turn requires checking all the assumptions of Theorem 2.1. We have
@@
Let $\beta,n$ be as in Theorem 1.2, and fix $t\in T$. Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2). Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$. Moreover, conditions 1, 2, 4 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem 2.1 applies.
We defer the proof to Section A.4. With Proposition 4.1, we deduce
@@
With the notation as in Theorem 2.1,
@@
Proof of Proposition 1.6.
Combining Lemmas 2.4, 3.2 and 3.5 gives
@@
$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(n\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\psi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\quad+\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{}\left[q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right]\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t$
@@
for $t\in T$, as the first summand is $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ while the last two are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$. Replacing $T$ by $T^{\prime}=[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}]$, the first summand is $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$ by Lemma 2.4. By positivity of the integrand, we bound the last two integrals over $T^{\prime}$ by those over $T$, which are themselves $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$. Now, all three summands are $o\mathopen{}\mathclose{{}\left(\sqrt{\beta\log\beta}}\right)$, as desired in Proposition 1.6. ∎
@@
4.2 Proof of Proposition 1.7
In this section, we prove Proposition 1.7.
@@
For any $a>0$ and $X_{1},\dots,X_{n}\in\mathbb{R}$, the number of modes of $\widehat{P}_{n}$ in $(a,\infty)$ is at most $|I|$ where $I=\{i\in[n]:X_{i}\geqslant a\}$.
Proof of Lemma 4.3.
@@
For $i\not\in I$, $g_{i}$ is monotonically decreasing on $[X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$. To this Gaussian mixture, we add in $g_{i}(t)$ for $i\in I$ one-by-one. By [CPW03, Theorem 2], each time the number of modes in $(a,\infty)$ increases by at most one. In $|I|$-many steps, there are at most $|I|$ such modes. ∎
@@
This argument crucially relies on the KDE being Gaussian. As discussed in [CPW03], the Gaussian kernel is the only kernel where for any fixed samples the number of modes of the KDE is non-increasing in the bandwidth $h$. For other kernels, we do suspect the analog of Lemma 4.3 to hold, but a different argument is needed. In particular, [MMF92, Mam95, KM97] avoids this problem by counting modes on compact sets.
Proof of Proposition 1.7.
By Lemma 4.3, symmetry of $T$ in 1.2 around $t=0$, linearity of expectations, and the tail bound $\mathbb{P}(|X|\geqslant a)\leqslant 2e^{-a^{2}/2}$ for $X\sim N(0,1)$
@@
by the definition of $\omega(\beta)$, proving Proposition 1.7. ∎
This concludes the proof of Theorem 1.2.
5 Concluding remarks
The question in the higher-dimensional case, as well as on the unit sphere $\mathbb{S}^{d-1}$ with uniformly distributed samples, remains open.
Acknowledgments
The authors would like to thank Enno Mammen for useful discussion and sharing important references. We also thank Dan Mikulincer for discussions on Gaussian approximation using Edgeworth expansions, Valeria Banica for comments on the method of stationary phase, and Alexander Zimin for providing Figure 2.
Funding. P.R. was supported by NSF grants DMS-2022448, CCF-2106377, and a gift from Apple. Y.S. was supported by the MIT UROP and MISTI France Programs.
Appendix A Additional proofs
A.1 Proof of 2.2
In this section we compute the first two moments of $(G,G^{\prime})$ to prove 2.2. Note that if $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, and for $t\in T$, then we have $\exp\Theta(t^{2}/\beta)\to 1$. This implies that exponentials in the moments are asymptotically $e^{-{t^{2}}/{2}}$. We frequently use the following fact about Gaussian integrals both in exact and asymptotic forms.
Fact A.1.

### v3

@@ 3.1 Bounding the third order error for small $y$
$\kappa_{t}^{\alpha}=\mathbb{E}\mathopen{}\mathclose{{\left[H^{\alpha}(Y(t))}}\right]=\mathbb{E}_{Z\sim N(0,I_{2})}\mathopen{}\mathclose{{\left[\frac{q_{t}(Z)}{\varphi(Z)}H^{\alpha}(Z)}}\right].$
If $|\alpha|\coloneqq\alpha_{1}+\alpha_{2}=s$, then $\kappa_{t}^{\alpha}$ can be bounded above asymptotically by the $s$-th moments of $\|Y\|$, which we bound in Section˜5.2.
@@
For $s\geqslant 3$, the cumulants of $Y$ with order $|\alpha|=s$ satisfy
$\kappa_{t}^{\alpha}\lesssim\eta_{s}\quad\text{where}\quad\eta_{s}\coloneqq\mathbb{E}[\|Y\|^{s}]\lesssim\mathopen{}\mathclose{{\left(\beta e^{t^{2}}}}\right)^{\frac{s-2}{4}}.$
Trivially, $|H^{(k,3-k)}(\mathbf{x})|\lesssim\|\mathbf{x}\|^{3}$. By Lemma˜2.3, we know for $y\geqslant\Delta_{t}\coloneqq{\delta_{t}+\sqrt{A_{t}/\alpha_{t}}}$ that
$\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})\asymp\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|$
Therefore, for any $t\in T$ and $y\geqslant\Delta_{t}$, we can bound
$\mathopen{}\mathclose{{\left|H^{(k,3-k)}\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)}}\right|\lesssim\|\tilde{y}\|^{3}.$ (3.4)
Now, by a similar method as Lemma˜2.3, we obtain the following bound. It says that when we integrate the Edgeworth series ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y\geqslant\Delta_{t}$ and $t\in T$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
Lemma $3.$
Recall $T,T^{\prime}$ from ˜1.2 and 1.3, and $A_{t}$ from Lemma˜2.3. Let $\Delta_{t}\coloneqq\delta_{t}+\sqrt{A_{t}/\alpha_{t}}$. Then
$\int_{T}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta},$
$\int_{T^{\prime}}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim\sqrt{\beta}.$
Proof of Lemma˜3.3.
Note that $y\geqslant\Delta_{t}$ if and only if $\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})\geqslant\sqrt{A_{t}}$. By Lemmas˜2.3 and 3.2, we can combine bounds ˜3.3 and 3.4 to obtain
$\int_{T}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim\sum_{k=0}^{3}\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\kappa_{t}^{(k,3-k)}\int_{\Delta_{t}}^{\infty}y\mathopen{}\mathclose{{\left[\varphi H^{(k,3-k)}}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\eta_{3}e^{-A_{t}}\int_{\Delta_{t}}^{\infty}ye^{-\frac{\tilde{y}^{2}}{2}}|\tilde{y}|^{3}\,\mathrm{d}y\,\mathrm{d}t$
$\asymp\int_{T}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}e^{-A_{t}}\eta_{3}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{\sqrt{A_{t}}}^{\infty}\tilde{y}^{4}e^{-\frac{\tilde{y}^{2}}{2}}d\tilde{y}}}\right)\,\mathrm{d}t$
@@
where we apply ˜2.11. The second statement for $T^{\prime}$ holds similarly by considering $\sup_{t\in T^{\prime}}\mathopen{}\mathclose{{\left(e^{\frac{t^{2}}{4}}}}\right)$. ∎
@@
One can see at this is actually an asymptotic equality by checking the Gaussian integrals in the proof above are of their typical order (i.e. no cancellation of leading terms). Hence, the decay is only a factor of $e^{-\omega(\beta)/4}$. For $t\not\in T$, even $t=\sqrt{2\log n-0.99\log\beta}$, the last inequality in Lemma˜3.3 fails and we get a bound of polynomially larger that $\sqrt{\beta}$. Then, as the third order error is asymptotically larger than the contribution of the Gaussian approximation, so the normal approximation is no longer valid. This can be seen by comparing the plots in Figure˜4.
@@ 3.2 Bounding higher order errors
We follow the classical proof of Theorem˜3.1 about the validity of the Edgeworth expansion as an asymptotic series to show bound the higher order pointwise error of density function as follows.
@@
Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$. Let $g_{2}\coloneqq q_{t}-\varphi$ and $g_{3}\coloneqq q_{t}-\varphi-n^{-\frac{1}{2}}\psi$. Then
$\sup_{\mathbf{x}\in\mathbb{R}^{2}}\mathopen{}\mathclose{{\left(1+\|\mathbf{x}\|^{s}}}\right)\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|(\mathbf{x})\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}$ (3.5)
as $n,\beta\to\infty$, for both $s=2$ and $s=3$, where we recall $\eta_{s}$ from Lemma˜3.2.
We defer the proof to Section˜5.4. Here, we comment on the differences with Theorem˜3.1: there, it is shown that order $s$ error is at most $n^{-\frac{s-1}{2}}$ provided $\eta_{s+1}$ is bounded. In our case, we pick up an extra $\eta_{s+1}$ factor as it is dependent on $\beta$. By Lemma˜3.2, this bound is at most
$n^{-\frac{s-1}{2}}\eta_{s+1}\lesssim\mathopen{}\mathclose{{\left(n^{-1}\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}}\right)^{\frac{s-1}{2}}\lesssim\begin{cases}e^{-\frac{(s-1)\omega(\beta)}{4}}&\textup{if }t\in T\\ \beta^{-(s-1)/2}&\textup{if }t\in T^{\prime}\end{cases}$ (3.6)
by definition of $T$ and $T^{\prime}$. This is the key reason for the extra $\omega(\beta)$ term in $T$: we get a small by non-negligible decay rate that matches Lemma˜3.3. Using Lemma˜3.5, we control the Kac-Rice integrals.
@@
Let $\Delta_{t}\coloneqq\delta_{t}+\sqrt{A_{t}/\alpha_{t}}$. If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for $c>0$, then asymptotically in $n,\beta\to\infty$
$\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
$\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$ $\lesssim e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta\log\beta}.$
Moreover, the left hand sides of both displays above with $T$ replaced by $T^{\prime}$ are $O(\sqrt{\beta})$.
Proof of Corollary˜3.6.
By Lemmas˜3.5 and 2.3, we let $\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})$ as before to obtain
$\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left|q_{t}-\varphi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\mathopen{}\mathclose{{\left(\frac{y}{1+A_{t}+\tilde{y}^{2}}}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\asymp e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{-\sqrt{A_{t}}}^{\sqrt{A_{t}}}\frac{|\tilde{y}|}{1+A_{t}+\tilde{y}^{2}}\,\mathrm{d}\tilde{y}}}\right)\,\mathrm{d}t$
$\asymp e^{-\frac{\omega(\beta)}{4}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\log\mathopen{}\mathclose{{\left(\frac{1+2A_{t}}{1+A_{t}}}}\right)\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta}\int_{T}\log(2)dt$
$\lesssim e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}$
where we check that if $y=0$, then $\tilde{y}=-\alpha_{t}^{\frac{1}{2}}\delta_{t}$ and $\alpha_{t}^{\frac{1}{2}}\delta_{t}\ll A_{t}^{\frac{1}{2}}$, so we may symmetrize the integral over $\tilde{y}$ up to a constant factor. Now similarly for the second display
@@
$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{3}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\int_{0}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left(A_{t}+\tilde{y}^{2}}}\right)^{\frac{3}{2}}}}\right)^{-1}\,\mathrm{d}y\,\mathrm{d}t$
$\lesssim e^{-\frac{\omega(\beta)}{2}}\int_{T}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\alpha_{t}^{-1}\mathopen{}\mathclose{{\left(\int_{-\infty}^{\infty}\frac{|\tilde{y}|}{1+|\tilde{y}|^{3}}\,\mathrm{d}\tilde{y}}}\right)\,\mathrm{d}t$
$\asymp e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta}\int_{T}\frac{2\pi}{3\sqrt{3}}dt$
$\asymp e^{-\frac{\omega(\beta)}{2}}\sqrt{\beta\log\beta}$
Now, the exact same computation but with $T$ replaced by $T^{\prime}$ gives bounds $O(\sqrt{\beta})$ upon replacing exponential in $\omega(\beta)$ decay rates with those in ˜3.6 for Lemma˜3.5.
In particular, by non-negativity of the integrand, the second equation holds upon replacing the bounds of integration of $y$ from $y\geqslant 0$ to $y\geqslant\Delta_{t}$. This is the form we will use.
4 Proof of Theorem˜1.2
We prove Propositions˜1.7 and 1.8 by checking ˜2.3, thereby proving Theorem˜1.2.
4.1 Proof of Proposition˜1.7
To prove Proposition˜1.7 we seek to apply Theorem˜2.1 to $\mathbb{E}U_{0}(F_{n},T)$. This in turn requires checking all the assumptions of Theorem˜2.1. We have
@@
Fix any $\beta>0$, an integer $n\geqslant 5$, and $t\in T$. Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2). Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$. Moreover, conditions 1, 2, 4 in Theorem˜2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem˜2.1 applies.
We defer the proof to Section˜5.5. With Proposition˜4.1, we deduce
@@
With the notation as in Theorem˜2.1,
@@
Proof of Proposition˜1.7.
We decompose the Kac-Rice integral as follows:
@@
$\quad+\int_{T}\int_{0}^{\Delta_{t}}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left[q_{t}-\varphi}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\quad+\int_{T}\int_{\Delta_{t}}^{\infty}\mathopen{}\mathclose{{\left(n\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\psi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
$\quad+\int_{T}\int_{\Delta_{t}}^{\infty}\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}y\mathopen{}\mathclose{{\left[q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right]\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t$
@@
Now, by Lemma˜2.4, the first summand is $\Theta\mathopen{}\mathclose{{\left(\sqrt{\beta\log\beta}}}\right)$, while the last three are $O\mathopen{}\mathclose{{\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}}\right)$ by Lemmas˜3.3 and 3.6. Replacing $T$ by $T^{\prime}$, all four summands are $O(\sqrt{\beta})$, as desired.
@@
4.2 Proof of Proposition˜1.8
In this section, we prove Proposition˜1.8.
@@
For any $a>0$ and $X_{1},\dots,X_{n}\in\mathbb{R}$, the number of modes of $\widehat{P}_{n}$ in $(a,\infty)$ is at most $|I|$ where $I=\{i\in[n]:X_{i}\geqslant a\}$. By symmetry, the same estimate holds for modes in $(-\infty,-a)$.
Proof of Lemma˜4.3.
@@
For $i\not\in I$, $g_{i}$ is monotonically decreasing on $X_{i},\infty)\supset(a,\infty)$, so $\sum_{i\not\in I}g_{i}(t)$ has no modes in $(a,\infty)$. To this Gaussian mixture, we add in $g_{i}(t)$ for $i\in I$ one-by-one. By [carreira2003number, Theorem 2], each time the number of modes in $(a,\infty)$ increases by at most one. In $|I|$-many steps, there are at most $|I|$ such modes. ∎
@@
Two remarks are in order:
  • As discussed in [carreira2003number], the scale-space property of the Gaussians allow us to view adding a component to the Gaussian mixture as adding a delta distribution to the mixture, which adds one mode, and applying a Gaussian blurring that does not create new modes.
  • This argument crucially relies on the KDE being Gaussian: as discussed in [carreira2003number], the Gaussian kernel is the only kernel where for any fixed samples the number of modes of the KDE is non-increasing in the bandwidth $h$, which enables the blurring step. For other kernels, we do suspect the analog of Lemma˜4.3 to hold, but a different argument is needed. In particular, [mammen91, mammen95, mammen97] avoids this problem by counting modes on compact sets.
Proof of Proposition˜1.8.
By Lemma˜4.3, symmetry of $T$ in ˜1.2 around $t=0$, linearity of expectations, and the tail bound $\mathbb{P}(|X|\geqslant a)\leqslant 2e^{-a^{2}/2}$ for $X\sim N(0,1)$
@@
by the definition of $\omega(\beta)$, proving Proposition˜1.8. ∎
Having proven Propositions˜1.7 and 1.8, we conclude Theorem˜1.2.
5 Additional proofs
5.1 Proof of Lemma˜5.1
We frequently make use of the following standard exercise on Gaussian integrals and dominated convergence.
Lemma 5.1.